# Visual Web Agents with Realistic Tasks

### v1

@@
(January 5, 2026)
[Uncaptioned image] WebGym: Scaling Training Environments for
Visual Web Agents with Realistic Tasks
@@ 1 Introduction
Vision-language models (VLMs) have recently demonstrated remarkable performance on agentic tasks on the web. These tasks often resemble those that a human user would aim to perform on the web, such as “Compare the prices and specs between AirPod 3 and AirPod 2 on Apple.com”. To solve such tasks, an agent will perform a series of actions, such as navigating to the relevant website, searching for both products, and finally, deciding on and submitting an answer to complete the task. Some web agents rely purely on textual information and try to solve tasks by simply utilizing accessibility trees webarena as the observation for the agent, while we focus on visual web agents, where agents observe screenshots, much like humans. This choice is based on the premise that web interaction semantics are defined by the rendered interface users see, and that reasoning over visual affordances enables more robust generalization across diverse and dynamically generated web pages (visual-web-arena; pae-webvoyager; test-time-interaction).
@@
There is a growing body of work on evaluating visual web agents by using benchmarks that involve either artificial or real-world websites. Benchmarks with artificial websites typically provide source code for hosting simulated environments that resemble real websites, such as REAL Bench (real-bench) and VisualWebArena (visual-web-arena). In contrast, realistic benchmarks such as GAIA-Web (gaia), BrowseComp (browsecomp), and InSTA-v3 (insta-v3) use tasks based on live websites. These “live-website” benchmarks tend to be more challenging, as real websites are inherently non-stationary (digirl; digi-q). For example, each time an agent visits the homepage of an e-commerce site, the set of displayed products may differ, and the same sequence of past actions could yield very different outcomes.
@@
Recent work has demonstrated that vision-language models can be adapted into web agents via supervised fine-tuning and task-specific training pipelines (glm-4.1v). However, in contrast to the rapid progress of online reinforcement learning (RL) in text-only domains such as software engineering and mathematical reasoning (swe-gym; e3), scaling RL for visual web agents remains substantially more challenging. Unlike text-based settings, where rollouts are fast and rewards are easily verifiable, visual web tasks require reasoning over fine-grained rendered interfaces and often lack unambiguous success signals (vision-r1; sarch2025grounded; online-mind2web), making data collection slow and inefficient. These challenges have inhibited scaling up post-training methods for visual web agents, particularly for diverse and long-horizon tasks, beyond current short-horizon settings (pae-webvoyager).
@@
Thus, to train advanced visual web agents, we first need a strong task set that includes problems which are challenging for existing models, and that span diverse domains and goals to encourage generalizable policy learning. A good task set should be able to support scaling performance by scaling these dimensions. These tasks must also be paired with clear evaluation protocols that provide meaningful learning signals. Such a task set would need to be substantially larger and more varied than existing ones (pae-webvoyager). In addition, utilizing a large task set requires a high training speed, with rollout efficiency being the main bottleneck. Rollouts in visual web environments typically take far longer than the learning updates themselves, largely due to the high computational overhead of simulating web browsers in current RL pipelines.
@@ 2 Related Work
Visual web agents for realistic tasks. This work focuses on visual web agents designed to solve real-world tasks using vision-language models (VLMs). Such agents emulate how humans interact with browsers on websites: they take website screenshots and metadata as input, and generate interactive actions supported by the browser (e.g., clicking or typing) to make progress toward a given goal. This process repeats until the agent outputs an action indicating task completion, resulting in multi-step trajectories. Prior to this work, both general-purpose VLMs and web-specific visual agents have shown promising capabilities on this problem of training visual web agents. General-purpose VLMs can perform reasonable grounding and reasoning on web tasks when combined with prompt engineering or light post-training on web-related data to improve general agent behavior, as demonstrated by models such as Gemma3 (gemma3), GLM-4.1v (glm-4.1v), and Qwen3-VL (qwen3-vl). In contrast, web-specific visual agents are models that are heavily post-trained on web/web-adjacent tasks, like UI-TARS (ui-tars).
@@
Evaluation benchmarks and training environments for realistic web agents. Recent work has made substantial progress in developing robust evaluation benchmarks for realistic web-browsing agents, while training environments remain relatively scarce. Existing benchmarks typically feature small, curated task sets with evaluation hints to assess specific reasoning or interaction abilities. Representative visual web benchmarks include WebVoyager (webvoyager), Mind2Web-Live (mind2web-live), and Mind2Web-2 (mind2web-2), alongside broader but non-visual ones such as AgentSynth (agentsynth), InSTA (insta-v3), BrowseComp (browsecomp), TravelPlanner (travelplanner), DeepShop (deepshop), and GAIA (gaia). A training environment, in contrast, must provide not only large-scale task sets but also efficient rollout and training systems. Efforts like WebRL (webrl) and the Proposer-Agent-Evaluator (PAE) framework (pae-webvoyager) have taken steps in this direction, but their task sets remain limited in diversity and hardness [^1] In this paper, “difficulty” or “difficulty level” specifically means the difficulty levels we propose as an attribute of the task set we propose, while we use the word “hardness” for indicating whether a given task is difficult or not., and their rollout systems are not optimized for speed. WebGym addresses these issues by expanding task diversity via rubric-derived fact groups and controlled decomposition into atomic and compositional variants (while limiting additional synthesis for sources like PAE-WebVoyager and AgentSynth to avoid near-duplicate synthetic goals), and by using a fully asynchronous rollout system that accelerates data collection and enables scalable reinforcement learning for web agents.
@@
Scaling online reinforcement learning for training agents. Prior work has improved the performance of LLMs and VLMs by scaling the size and diversity of supervised fine-tuning (SFT) datasets, such as Super Natural-Instruct (super-natural-instruct) and OpenThoughts (open-thoughts). More recently, studies have shown that increasing the number of training trajectories through online RL can further boost agent capabilities, even with algorithms like REINFORCE (original-reinforce), across software engineering (swe-gym), math reasoning (e3), deep research (deepresearcher-rl; search-r1; paper-search-agent), and general reasoning. WebGym builds on these insights by constructing a large, diverse task set spanning multiple domains and difficulty levels, and training agents via on-policy RL from collected rollouts.
@@ 3 Tasks in WebGym and Evaluation Protocol
Training with online RL requires a large number of model-generated rollouts guided by reliable reward signals on a family of tasks (“environments” in traditional multi-task RL nomenclature) that we curate. To learn generalizable policies with such a procedure, the task set must (1) span diverse domains and websites, (2) cover a range of “difficulties” from atomic to compositional tasks, and (3) include verifiable evaluators that translate outcomes into meaningful learning signals. Existing web agent benchmarks, while effective for evaluation, are typically too small and static for large-scale training. Existing training tasks are either non-diverse (pae-webvoyager) and lack a wide range of task difficulties (insta-v3), or do not come paired with good task-level evaluation criteria, rubrics, or guidance for enabling robust policy evaluation (insta-v3; pae-webvoyager; agentsynth; browsecomp; gaia). We describe how we procedurally construct a task set that satisfies these desiderata, with assistance from LLMs.
@@
InSTA-v3 (insta-v3) 146,348 146,441
PAE-WebVoyager (pae-webvoyager) 13 128,499
AgentSynth-Web (agentsynth) 328 2,086
BrowseComp (browsecomp) 7 1,266
TravelPlanner (travelplanner) 1 1,225
Mind2Web-Live (mind2web-live) 76 542
Online Mind2Web (online-mind2web) 139 300
DeepShop (deepshop) 1 150
Mind2Web-2 (mind2web-2) 44 130
GAIA-Web (gaia) 1 87
@@ 3.1 Obtaining a Seed Set of Tasks
To obtain a set of seed tasks, we collect 10 of the most recent and highly utilized task sets in the open-source community, as listed in Table 1. These tasks come from both existing benchmarks and existing training sets. Importantly, we do not designate any entire benchmark as “held out”, because many benchmarks share websites, domains, or task patterns with one another, making train-test separation at the level of existing benchmarks unreliable. Instead, WebGym constructs its own train-test split at the task and website level to ensure strict separation without discarding large amounts of useful data. Specifically, our tasks come from a) InSTA-v3 (insta-v3), which provides wide coverage of websites and domains; b) PAE-WebVoyager (pae-webvoyager) and AgentSynth (agentsynth), which provide synthetic tasks focused on narrow domains; c) BrowseComp (browsecomp) and GAIA-Web (gaia), which include difficult tasks with concrete, verifiable answers; d) TravelPlanner (travelplanner) and DeepShop (deepshop), which focus on domain-specific browsing problems; and e) Mind2Web-Live (mind2web-live), Online Mind2Web (online-mind2web), and Mind2Web-2 (mind2web-2), which contain human-verified high-quality tasks that are indeed solvable. We next describe how we construct our full task set.
@@ 3.2 Task Set Construction Procedure
Recent research has demonstrated strong improvements when constructing curriculum on tasks with similar goals but different hardness (test-time-interaction; e3). In parallel, other efforts show that developing evaluation rubrics can significantly improve evaluation accuracy (task-specific-rubric; rubric-anchors; rubrics-as-rewards). We combine these ideas by generating evaluation rubrics and using them to construct decomposed tasks. We elaborate our task construction algorithm in Algorithm 1 and an example in Figure 2. All constructions are generated using GPT-4o.
@@
As a result, tasks generated by WebGym (1) avoid being trivial by restricting to sufficiently large groups, (2) are strictly easier than the original task, yielding denser reward signals, and (3) are guaranteed to be well-defined whenever the original task is well-defined. In contrast, PAE (pae-webvoyager) generates synthetic tasks that may be ill-posed or impossible to solve. While WebGym is related to the approach in AgentSynth (agentsynth), it does not require executing agent rollouts to generate new tasks making it cheaper.
@@ 3.4 Statistics: Size, Breadth and Depth
Figure 4: Analysis of the WebGym task set. Left: distribution of tasks across domains according to the Mind2Web-2 taxonomy (mind2web-2). Right: Distribution of trajectory lengths by task difficulty for answered trajectories over multiple iterations. For trajectories working on medium- and hard-difficulty tasks that exceeds 30 tasks during evaluation, we filter in only trajectories under 30 tasks to make the comparison fair. The violin plots show that the probability density of trajectory lengths, with the KDE mode (black line, KDE mode is the peak of the smoothed probability density curve, where the violin is widest) indicating the most likely length and mean (red) indicating the average.
@@ 3.5 Evaluator and Reward Function Design
Evaluating long-horizon visual agents is challenging because the produced trajectories often make partial progress, end up at visually similar but incorrect pages, and generate answers that may appear correct unless verified against grounded evidence. Unlike math reasoning or coding, “reference” answers of any sort are rare to obtain for most task instructions because only a small minority of benchmarks provide them (for example BrowseComp and GAIA-Web (browsecomp; gaia)) and the notion of a reference answer does not exist in many cases (e.g., tasks that do not involve question answering). This difference reflects a fundamental challenge in designing open-ended web tasks with a binary answer. Thus, evaluation cannot rely on direct matching to ground-truth outputs and must instead interpret the trajectory in context. To this end, we mainly focus on LLM-judged evaluation, while reference answers, when available in the source benchmark, overrides the judge.
For faithful evaluation, recent research has adopted a criterion proposal approach that generates task-specific criteria to guide evaluation (k2; chen2025rm). However, we posit that tasks usually come with structures, so we propose to generate rubric with structures such as fact groups. As discussed in Section 3.2, in WebGym, each task is paired with a rubric of several fact groups, where each group contains one or more criteria that guides evaluation reducing false positives from ambiguity about task success.
With the task-specific rubric, a trajectory receives reward only when all criteria are satisfied, producing a binary signal consistent across domains and difficulty levels. Since not all screenshots are informative, the judge focuses on evidence-bearing pages, obtained via a high-recall keypoint selection process inspired by Online Mind2Web (online-mind2web) that retains relevant observations while filtering out distractors such as detours and ads. In practice, we use GPT-4o for both keypoint selection (prompt shown in §E.2 and per-criterion evaluation (prompt shown in  §E.3). The evaluator proposes keypoint screenshots from the task description, iterates through trajectory steps, and issues binary judgments for each criterion using the selected keypoints and final answer.
@@ 4 Rollout System in WebGym
To our knowledge, the WebGym provides the the first, open-source rollout system optimized for asynchronous multi-turn simulations for web agent tasks. In contrast to work on RL infrastructure for LLMs, WebGym focuses on an orthogonal dimension: speeding up rollout generation up by introducing a highly efficient simulation server for web browsing so that the rollout pipeline does not synchronize between any step or episode. This server can be easily integrated with any RL framework that focuses on system-level speedup like optimizing the weights save/load overhead and algorithmic speedup like introducing episode/step/token-level off-policy RL. Examples of these RL frameworks include veRL (verl), veRL-agent (verl-agent), PipelineRL (pipeline-rl), SkyRL (skyrl-agent), and AReaL (areal).

### v2

@@
(January 7, 2026)
@@ 1 Introduction
Vision-language models (VLMs) have recently demonstrated remarkable performance on agentic tasks on the web. These tasks often resemble those that a human user would aim to perform on the web, such as “Compare the prices and specs between AirPod 3 and AirPod 2 on Apple.com”. To solve such tasks, an agent will perform a series of actions, such as navigating to the relevant website, searching for both products, and finally, deciding on and submitting an answer to complete the task. Some web agents rely purely on textual information and try to solve tasks by simply utilizing accessibility trees Zhou et al. (2023) as the observation for the agent, while we focus on visual web agents, where agents observe screenshots, much like humans. This choice is based on the premise that web interaction semantics are defined by the rendered interface users see, and that reasoning over visual affordances enables more robust generalization across diverse and dynamically generated web pages (Koh et al., 2024; Zhou et al., 2024; Shen et al., 2025).
@@
There is a growing body of work on evaluating visual web agents by using benchmarks that involve either artificial or real-world websites. Benchmarks with artificial websites typically provide source code for hosting simulated environments that resemble real websites, such as REAL Bench (Garg et al., 2025) and VisualWebArena (Koh et al., 2024). In contrast, realistic benchmarks such as GAIA-Web (Mialon et al., 2023), BrowseComp (Wei et al., 2025), and InSTA-v3 (Trabucco et al., 2025) use tasks based on live websites. These “live-website” benchmarks tend to be more challenging, as real websites are inherently non-stationary (Bai et al., 2024, 2025a). For example, each time an agent visits the homepage of an e-commerce site, the set of displayed products may differ, and the same sequence of past actions could yield very different outcomes.
@@
Recent work has demonstrated that vision-language models can be adapted into web agents via supervised fine-tuning and task-specific training pipelines (Hong et al., 2025). However, in contrast to the rapid progress of online reinforcement learning (RL) in text-only domains such as software engineering and mathematical reasoning (Pan et al., 2024a; Setlur et al., 2025), scaling RL for visual web agents remains substantially more challenging. Unlike text-based settings, where rollouts are fast and rewards are easily verifiable, visual web tasks require reasoning over fine-grained rendered interfaces and often lack unambiguous success signals (Huang et al., 2025a; Sarch et al., 2025; Xue et al., 2025), making data collection slow and inefficient. These challenges have inhibited scaling up post-training methods for visual web agents, particularly for diverse and long-horizon tasks, beyond current short-horizon settings (Zhou et al., 2024).
@@
Thus, to train advanced visual web agents, we first need a strong task set that includes problems which are challenging for existing models, and that span diverse domains and goals to encourage generalizable policy learning. A good task set should be able to support scaling performance by scaling these dimensions. These tasks must also be paired with clear evaluation protocols that provide meaningful learning signals. Such a task set would need to be substantially larger and more varied than existing ones (Zhou et al., 2024). In addition, utilizing a large task set requires a high training speed, with rollout efficiency being the main bottleneck. Rollouts in visual web environments typically take far longer than the learning updates themselves, largely due to the high computational overhead of simulating web browsers in current RL pipelines.
@@ 2 Related Work
Visual web agents for realistic tasks. This work focuses on visual web agents designed to solve real-world tasks using vision-language models (VLMs). Such agents emulate how humans interact with browsers on websites: they take website screenshots and metadata as input, and generate interactive actions supported by the browser (e.g., clicking or typing) to make progress toward a given goal. This process repeats until the agent outputs an action indicating task completion, resulting in multi-step trajectories. Prior to this work, both general-purpose VLMs and web-specific visual agents have shown promising capabilities on this problem of training visual web agents. General-purpose VLMs can perform reasonable grounding and reasoning on web tasks when combined with prompt engineering or light post-training on web-related data to improve general agent behavior, as demonstrated by models such as Gemma3 (Gemma-Team et al., 2025), GLM-4.1v (Hong et al., 2025), and Qwen3-VL (Bai et al., 2025b). In contrast, web-specific visual agents are models that are heavily post-trained on web/web-adjacent tasks, like UI-TARS (Qin et al., 2025).
@@
Evaluation benchmarks and training environments for realistic web agents. Recent work has made substantial progress in developing robust evaluation benchmarks for realistic web-browsing agents, while training environments remain relatively scarce. Existing benchmarks typically feature small, curated task sets with evaluation hints to assess specific reasoning or interaction abilities. Representative visual web benchmarks include WebVoyager (He et al., 2024), Mind2Web-Live (Pan et al., 2024b), and Mind2Web-2 (Gou et al., 2025), alongside broader but non-visual ones such as AgentSynth (Xie et al., 2025), InSTA (Trabucco et al., 2025), BrowseComp (Wei et al., 2025), TravelPlanner (Xie et al., 2024), DeepShop (Lyu et al., 2025), and GAIA (Mialon et al., 2023). A training environment, in contrast, must provide not only large-scale task sets but also efficient rollout and training systems. Efforts like WebRL (Qi et al., 2024) and the Proposer-Agent-Evaluator (PAE) framework (Zhou et al., 2024) have taken steps in this direction, but their task sets remain limited in diversity and hardness [^1] In this paper, “difficulty” or “difficulty level” specifically means the difficulty levels we propose as an attribute of the task set we propose, while we use the word “hardness” for indicating whether a given task is difficult or not., and their rollout systems are not optimized for speed. WebGym addresses these issues by expanding task diversity via rubric-derived fact groups and controlled decomposition into atomic and compositional variants (while limiting additional synthesis for sources like PAE-WebVoyager and AgentSynth to avoid near-duplicate synthetic goals), and by using a fully asynchronous rollout system that accelerates data collection and enables scalable reinforcement learning for web agents.
@@
Scaling online reinforcement learning for training agents. Prior work has improved the performance of LLMs and VLMs by scaling the size and diversity of supervised fine-tuning (SFT) datasets, such as Super Natural-Instruct (Wang et al., 2022) and OpenThoughts (Guha et al., 2025). More recently, studies have shown that increasing the number of training trajectories through online RL can further boost agent capabilities, even with algorithms like REINFORCE (Williams, 1992), across software engineering (Pan et al., 2024a), math reasoning (Setlur et al., 2025), deep research (Zheng et al., 2025; Jin et al., 2025; He et al., 2025), and general reasoning. WebGym builds on these insights by constructing a large, diverse task set spanning multiple domains and difficulty levels, and training agents via on-policy RL from collected rollouts.
@@ 3 Tasks in WebGym and Evaluation Protocol
Training with online RL requires a large number of model-generated rollouts guided by reliable reward signals on a family of tasks (“environments” in traditional multi-task RL nomenclature) that we curate. To learn generalizable policies with such a procedure, the task set must (1) span diverse domains and websites, (2) cover a range of “difficulties” from atomic to compositional tasks, and (3) include verifiable evaluators that translate outcomes into meaningful learning signals. Existing web agent benchmarks, while effective for evaluation, are typically too small and static for large-scale training. Existing training tasks are either non-diverse (Zhou et al., 2024) and lack a wide range of task difficulties (Trabucco et al., 2025), or do not come paired with good task-level evaluation criteria, rubrics, or guidance for enabling robust policy evaluation (Trabucco et al., 2025; Zhou et al., 2024; Xie et al., 2025; Wei et al., 2025; Mialon et al., 2023). We describe how we procedurally construct a task set that satisfies these desiderata, with assistance from LLMs.
@@
InSTA-v3 (Trabucco et al., 2025) 146,348 146,441
PAE-WebVoyager (Zhou et al., 2024) 13 128,499
AgentSynth-Web (Xie et al., 2025) 328 2,086
BrowseComp (Wei et al., 2025) 7 1,266
TravelPlanner (Xie et al., 2024) 1 1,225
Mind2Web-Live (Pan et al., 2024b) 76 542
Online Mind2Web (Xue et al., 2025) 139 300
DeepShop (Lyu et al., 2025) 1 150
Mind2Web-2 (Gou et al., 2025) 44 130
GAIA-Web (Mialon et al., 2023) 1 87
@@ 3.1 Obtaining a Seed Set of Tasks
To obtain a set of seed tasks, we collect 10 of the most recent and highly utilized task sets in the open-source community, as listed in Table 1. These tasks come from both existing benchmarks and existing training sets. Importantly, we do not designate any entire benchmark as “held out”, because many benchmarks share websites, domains, or task patterns with one another, making train-test separation at the level of existing benchmarks unreliable. Instead, WebGym constructs its own train-test split at the task and website level to ensure strict separation without discarding large amounts of useful data. Specifically, our tasks come from a) InSTA-v3 (Trabucco et al., 2025), which provides wide coverage of websites and domains; b) PAE-WebVoyager (Zhou et al., 2024) and AgentSynth (Xie et al., 2025), which provide synthetic tasks focused on narrow domains; c) BrowseComp (Wei et al., 2025) and GAIA-Web (Mialon et al., 2023), which include difficult tasks with concrete, verifiable answers; d) TravelPlanner (Xie et al., 2024) and DeepShop (Lyu et al., 2025), which focus on domain-specific browsing problems; and e) Mind2Web-Live (Pan et al., 2024b), Online Mind2Web (Xue et al., 2025), and Mind2Web-2 (Gou et al., 2025), which contain human-verified high-quality tasks that are indeed solvable. We next describe how we construct our full task set.
@@ 3.2 Task Set Construction Procedure
Recent research has demonstrated strong improvements when constructing curriculum on tasks with similar goals but different hardness (Shen et al., 2025; Setlur et al., 2025). In parallel, other efforts show that developing evaluation rubrics can significantly improve evaluation accuracy (Pathak et al., 2025; Huang et al., 2025b; Gunjal et al., 2025). We combine these ideas by generating evaluation rubrics and using them to construct decomposed tasks. We elaborate our task construction algorithm in Algorithm 1 and an example in Figure 2. All constructions are generated using GPT-4o.
@@
As a result, tasks generated by WebGym (1) avoid being trivial by restricting to sufficiently large groups, (2) are strictly easier than the original task, yielding denser reward signals, and (3) are guaranteed to be well-defined whenever the original task is well-defined. In contrast, PAE (Zhou et al., 2024) generates synthetic tasks that may be ill-posed or impossible to solve. While WebGym is related to the approach in AgentSynth (Xie et al., 2025), it does not require executing agent rollouts to generate new tasks making it cheaper.
@@ 3.4 Statistics: Size, Breadth and Depth
Figure 4: Analysis of the WebGym task set. Left: distribution of tasks across domains according to the Mind2Web-2 taxonomy (Gou et al., 2025). Right: Distribution of trajectory lengths by task difficulty for answered trajectories over multiple iterations. For trajectories working on medium- and hard-difficulty tasks that exceeds 30 tasks during evaluation, we filter in only trajectories under 30 tasks to make the comparison fair. The violin plots show that the probability density of trajectory lengths, with the KDE mode (black line, KDE mode is the peak of the smoothed probability density curve, where the violin is widest) indicating the most likely length and mean (red) indicating the average.
@@ 3.5 Evaluator and Reward Function Design
Evaluating long-horizon visual agents is challenging because the produced trajectories often make partial progress, end up at visually similar but incorrect pages, and generate answers that may appear correct unless verified against grounded evidence. Unlike math reasoning or coding, “reference” answers of any sort are rare to obtain for most task instructions because only a small minority of benchmarks provide them (for example BrowseComp and GAIA-Web (Wei et al., 2025; Mialon et al., 2023)) and the notion of a reference answer does not exist in many cases (e.g., tasks that do not involve question answering). This difference reflects a fundamental challenge in designing open-ended web tasks with a binary answer. Thus, evaluation cannot rely on direct matching to ground-truth outputs and must instead interpret the trajectory in context. To this end, we mainly focus on LLM-judged evaluation, while reference answers, when available in the source benchmark, overrides the judge.
For faithful evaluation, recent research has adopted a criterion proposal approach that generates task-specific criteria to guide evaluation (Kimi-Team et al., 2025; Chen et al., 2025). However, we posit that tasks usually come with structures, so we propose to generate rubric with structures such as fact groups. As discussed in Section 3.2, in WebGym, each task is paired with a rubric of several fact groups, where each group contains one or more criteria that guides evaluation reducing false positives from ambiguity about task success.
With the task-specific rubric, a trajectory receives reward only when all criteria are satisfied, producing a binary signal consistent across domains and difficulty levels. Since not all screenshots are informative, the judge focuses on evidence-bearing pages, obtained via a high-recall keypoint selection process inspired by Online Mind2Web (Xue et al., 2025) that retains relevant observations while filtering out distractors such as detours and ads. In practice, we use GPT-4o for both keypoint selection (prompt shown in §E.2 and per-criterion evaluation (prompt shown in  §E.3). The evaluator proposes keypoint screenshots from the task description, iterates through trajectory steps, and issues binary judgments for each criterion using the selected keypoints and final answer.
@@ 4 Rollout System in WebGym
To our knowledge, the WebGym provides the the first, open-source rollout system optimized for asynchronous multi-turn simulations for web agent tasks. In contrast to work on RL infrastructure for LLMs, WebGym focuses on an orthogonal dimension: speeding up rollout generation up by introducing a highly efficient simulation server for web browsing so that the rollout pipeline does not synchronize between any step or episode. This server can be easily integrated with any RL framework that focuses on system-level speedup like optimizing the weights save/load overhead and algorithmic speedup like introducing episode/step/token-level off-policy RL. Examples of these RL frameworks include veRL (Sheng et al., 2025), veRL-agent (Feng et al., 2025), PipelineRL (Piché et al., 2025), SkyRL (Cao et al., 2025), and AReaL (Fu et al., 2025).